# Leveraging (Biased) Information: Multi-armed Bandits with Offline Data

### v1

@@ 1.1 Related Works
Notation. We denote ${\cal N}(\mu,1)$ as the Gaussian disribution with mean $\mu$ and variance 1. We abbreviate identically and independently distributed as “i.i.d.”. The relationship $R\sim P$ means that the random variable $R$ follows the probability distribution $P$.
2 Model
@@
We consider a stochastic $K$-armed bandit model with possibly biased offline data. Denote ${\cal A}=\{1,\ldots,K\}$ as the set of $K$ arms. The event horizon consists of a warm-start phase, then an online phase. During the warm-start phase, the DM receive a collection of $T_{\text{S}}(a)\in\mathbb{Z}_{\geq 0}$ i.i.d. samples, denoted $S(a)=\{X_{s}(a)\}^{T_{\text{S}}(a)}_{s=1}$, for each $a\in{\cal A}$. We postulate that $X_{1}(a),\ldots,X_{T_{\text{S}}(a)}(a)\sim P^{\text{(off)}}_{a}$, the offline reward distribution on arm $a$.
@@
The subsequent online phase consists of $T$ time steps. For time steps $t=1,\ldots,T$, the DM chooses an arm $A_{t}\in{\cal A}$ to pull, which gives the DM a random reward $R_{t}$. Pulling arm $a$ in the online phase generates a random reward $R(a)\sim P^{\text{(on)}}_{a}$, dubbed the online reward distribution on arm $a$. We emphasize that $P^{\text{(on)}}=\{P^{\text{(on)}}_{a}\}_{a\in{\cal A}}$ needs not be the same as $P^{\text{(off)}}=\{P^{\text{(off)}}_{a}\}_{a\in{\cal A}}$. Finally, the DM proceeds to time step $t+1$, or terminates when $t=T$.
@@
The DM chooses arms $A_{1},\ldots,A_{T}$ with a non-anticipatory policy $\pi=\{\pi_{t}\}^{\infty}_{t=1}$. The function $\pi_{t}$ maps the observations $H_{t-1}=(S=\{S(a)\}_{a\in{\cal A}},\{(A_{s}.R_{s})\}^{t-1}_{s=1})$ collected at the end of $t-1$, to an element in $\{x\in\mathbb{R}^{K}_{\geq 0}:\sum_{a\in{\cal A}}x_{a}=1\}$. The quantity $\pi_{t}(a~{}|~{}H_{t-1})\in[0,1]$ is the probability of $A_{t}=a$, conditioned on $H_{t-1}$ under policy $\pi$. While our proposed algorithms are deterministic, i.e. $\pi_{t}(a~{}|~{}H_{t-1})\in\{0,1\}$ always, we allow randomized policies in our regret lower bound results. The policy $\pi$ could utilize $S$, which provides a potentially biased estimate to the online model since we allow $P^{\text{(on)}}\neq P^{\text{(off)}}$.
@@
Altogether, the underlying instance $I$ is specified by the tuple $({\cal A},\{T_{\text{S}}(a)\}_{a\in{\cal A}},P,T)$, where $P=(P^{\text{(off)}},P^{\text{(on)}})$. The DM only knows ${\cal A}$ and $S$, but not $P$ and $T$, at the start of the online phase. For every $a\in{\cal A}$, we assume that both $P^{\text{(off)}}_{a},P^{\text{(on)}}_{a}$ are 1-subGaussian, which is also known to the DM. For $a\in{\cal A}$, denote $\mu^{\text{(on)}}(a)=\mathbb{E}_{R(a)\sim P^{\text{(on)}}_{a}}[R(a)]$, and $\mu^{\text{(off)}}(a)=\mathbb{E}_{R(a)\sim P^{\text{(off)}}_{a}}[R(a)]$. We do not have any boundedness assumptions on $\mu^{\text{(on)}}(a),\mu^{\text{(off)}}(a)$.
@@
$\text{Reg}_{T}(\pi,P)=T\mu_{*}^{\text{(on)}}-\sum_{t=1}^{T}\mu^{\text{(on)}}(A_{t})=\sum^{T}_{t=1}\Delta(A_{t})$ (3)
@@
Auxiliary input: valid bias bound. We say that $V=\{V(a)\}_{a\in{K}}\in(\mathbb{R}\cup\{\infty\})^{|\mathcal{A}|}$ is a valid bias bound on an instance $I$ if
$V(a)\geq|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|\quad\text{for each $a\in{\cal A}$.}$ (4)
@@
In our forthcoming policy MIN-UCB, the DM is endowed the auxiliary input $V$ in addition to the offline dataset $S$ before the online phase begins. The quantity $V(a)$ serves as an upper bound on the amount of distributional shift from $P^{\text{(off)}}_{a}$ to $P^{\text{(on)}}_{a}$. The condition (4) always holds in the trivial case of $V(a)=\infty$, which is when the DM has no additional knowledge. By contrast, when $V(a)\neq\infty$, the DM has a non-trivial knowledge on the difference $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$.
@@
The knowledge on an upper bound such as $V$ is in line with the model assumptions in research works on learning under distributional shift. Similar upper bounds are assumed in the contexts of supervised learning (Crammer et al., 2008), stochastic optimization (Besbes et al., 2022), offline policy learning on contextual bandits (Si et al., 2023), multi-task bandit learning (Wang et al., 2021), for example. Methodologies for constructing such upper bounds are provided in some literature. For instance, (Blanchet et al., 2019) designs an approach based on several machine learning estimators, such as LASSO. (Si et al., 2023) provides some managerial insights on how to construct them empirically. (Chen et al., 2022) constructs them through cross-validation, and implement it in a case study.
@@ 3 An Impossibility Result
We illustrate the impossibility result with two instances $I_{P},I_{Q}$. The instances $I_{P},I_{Q}$ share the same ${\cal A}=\{1,2\},\{T_{\text{S}}(a)\}_{a\in{\cal A}}$ and $T$. However, $I_{P},I_{Q}$ differ in their respective reward distributions $P$, $Q$, and $I_{P},I_{Q}$ have different optimal arms. Consider a fixed but arbitrary $\beta\in(0,1/2)$. Instance $I_{P}$ has well-aligned reward distributions $P^{\text{(off)}}=P^{\text{(on)}}$, where
$P^{\text{(on)}}_{1}={\cal N}(0,1),\quad P^{\text{(on)}}_{2}={\cal N}(-T^{-\beta},1).$
Note that $\Delta(2)=T^{-\beta}$. In $I_{P}$, an offline dataset provides useful hints on arm 1 being optimal. For example, given $T_{\text{S}}(1)=T_{\text{S}}(2)\geq 128(T^{2\beta}-T^{2\beta-\epsilon})\log T$ where $\epsilon\in(0,\beta)$, the existing policies H-UCB and MonUCB achieve an expected regret (see Equation (2)) at most
$O\left(\frac{\log T}{\Delta(2)}-T_{\text{S}}(2)\Delta(2)\right)$ $=O(\log T(T^{\beta}-T^{\beta}+T^{\beta-\epsilon}))$
$=O(T^{\beta-\epsilon}\log T),$
which strictly outperforms the vanilla UCB policy that incurs expected regret $O(\log T/\Delta(2))=O(T^{\beta}\log T)$. Despite the apparent improvement, the following claim shows that any non-anticipatory policy (not just H-UCB or MonUCB) that outperforms the vanilla UCB on $I_{P}$ would incur a larger regret than the vanilla UCB on a suitably chosen $I_{Q}$.
Proposition 3.1.
Let $T_{\text{S}}(1),T_{\text{S}}(2)$ be arbitrary. Consider an arbitrary non-anticipatory policy $\pi$ (which only possesses the offline dataset $S$ but not the auxiliary input $V$) satisfies $\mathbb{E}[\text{Reg}_{T}(\pi,P)]\leq CT^{\beta-\epsilon}\log T$ on instance $I_{P}$, where $\epsilon\in(0,\beta)$, $C>0$ is an absolute constant, and the horizon $T$ is so large that $C<T^{\epsilon}/(4\log T)$. Set $Q=(Q^{\text{(off)}},Q^{\text{(on)}})$ in the instance $I_{Q}$ as
$Q^{\text{(off)}}$ $=P^{\text{(off)}},\quad Q^{\text{(on)}}_{1}=P^{\text{(on)}}_{1}\text{, but}$
@@
The following inequality holds:
$\mathbb{E}[\text{Reg}(\pi,Q)]$ $\geq\frac{T^{1-\beta}}{4}\cdot e^{-2}-CT^{\beta-\epsilon}\log T$ (5)
$=\Omega(T^{1-\beta})>\Omega(\sqrt{T}).$
@@
Proposition 3.1 is proved in Appendix section C.2. We first remark on $I_{Q}$. Different from $I_{P}$, arm 2 is the optimal arm in $I_{Q}$. Thus, the offline data from $Q^{\text{(off)}}$ provides the misleading information that arm 1 has a higher expected reward. In addition, note that $Q^{\text{(off)}}=P^{\text{(off)}}$, so the offline datasets in instances $I_{P},I_{Q}$ are identically distributed. The proposition suggests that no non-anticipatory policy can simultaneously (i) establish that $S$ is useful for online learning in $I_{P}$ to improve upon the UCB policy, (ii) establish that $S$ is to be disregarded for online learning in $I_{Q}$ to match the UCB policy, negating the short-lived conjecture in (Q).
@@
Lastly, we allow $T_{\text{S}}(1),T_{\text{S}}(2)$ to be arbitrary. In the most informative case of $T_{\text{S}}(1)=T_{\text{S}}(2)=\infty$, meaning that the DM knows $P^{\text{(off)}},Q^{\text{(off)}}$ in $I_{P},I_{Q}$ respectively, the DM still cannot achieve (i, ii) simultaneously. Indeed, the assumed improvement $\mathbb{E}[\text{Reg}_{T}(\pi,P)]\leq CT^{\beta-\epsilon}\log T$ limits the number of pulls on arm 2 in $I_{P}$ during the online phase. We harness this property to construct an upper bound on the KL divergence between on online dynamics on $I_{P}$, $I_{Q}$, which in turns implies the DM cannot differentiate between $I_{P},I_{Q}$ under policy $\pi$. The KL divergence argument utilizes a chain rule (see Appendix C.1) on KL divergence adapted to our setting.
4 Design and Analysis of MIN-UCB

### v2

@@ 1.4  Notation
We denote ${\cal N}(\mu,1)$ as the Gaussian distribution with mean $\mu$ and variance 1. We abbreviate “identically and independently distributed” as “i.i.d.”. The relationship $R\sim P$ means that the random variable $R$ follows the probability distribution $P$.
2  Model: Multi-armed Bandits with Offline Data
@@
We consider a stochastic $K$-armed bandit model with possibly biased offline data. Denote ${\cal A}=\{1,\ldots,K\}$ as the set of $K$ arms. The event horizon consists of a warm-start phase, followed by an online phase. During the warm-start phase, the DM receives a collection of $T_{\text{S}}(a)\in\mathbb{Z}_{\geq 0}$ i.i.d. samples, denoted $S(a)=\{X_{s}(a)\}^{T_{\text{S}}(a)}_{s=1}$, for each $a\in{\cal A}$. We denote the set of offline samples collectively as $S=\{S(a)\}_{a\in{\cal A}}$. We postulate that $X_{1}(a),\ldots,X_{T_{\text{S}}(a)}(a)\sim P^{\text{(off)}}_{a}$, the offline reward distribution on arm $a$.
@@
The subsequent online phase consists of $T$ time steps. For $t=1,\ldots,T$, the DM chooses an arm $A_{t}\in{\cal A}$ to pull, which gives the DM a random reward $R_{t}$. Pulling arm $a$ in the online phase generates a random reward $R(a)\sim P^{\text{(on)}}_{a}$, dubbed the online reward distribution on arm $a$. We emphasize that $P^{\text{(on)}}=\{P^{\text{(on)}}_{a}\}_{a\in{\cal A}}$ needs not be the same as $P^{\text{(off)}}=\{P^{\text{(off)}}_{a}\}_{a\in{\cal A}}$. Finally, the DM proceeds to time step $t+1$, or terminates when $t=T$.
@@
The DM chooses arms $A_{1},\ldots,A_{T}$ with a non-anticipatory policy $\pi=\{\pi_{t}\}^{\infty}_{t=1}$. The function $\pi_{t}$ maps the observations $H_{t-1}=(S,\{(A_{s}.R_{s})\}^{t-1}_{s=1})$ collected at the end of t1t-1, to an element in $\{x\in\mathbb{R}^{K}_{\geq 0}:\sum_{a\in{\cal A}}x_{a}=1\}$. The quantity $\pi_{t}(a~|~H_{t-1})\in[0,1]$ is the probability of $A_{t}=a$, conditioned on $H_{t-1}$ under policy $\pi$. While our proposed algorithms are deterministic, i.e. $\pi_{t}(a~|~H_{t-1})\in\{0,1\}$ always, we allow randomized policies in our regret lower bound results. The policy $\pi$ could utilize $S$, which provides a potentially biased estimate to the online model since we allow $P^{\text{(on)}}\neq P^{\text{(off)}}$.
@@
Altogether, the underlying instance $I$ is specified by the tuple $({\cal A},\{T_{\text{S}}(a)\}_{a\in{\cal A}},P,T)$, where $P=(P^{\text{(off)}},P^{\text{(on)}})$. The DM only knows $\cal A$} and $S$, but not $P$ and $T$, at the start of the online phase. For every $a\in{\cal A}$, we assume that both $P^{\text{(off)}}_{a},P^{\text{(on)}}_{a}$ are 1-subGaussian, which is also known to the DM. For $a\in{\cal A}$, denote $\mu^{\text{(on)}}(a)=\mathbb{E}_{R(a)\sim P^{\text{(on)}}_{a}}[R(a)]$, and $\mu^{\text{(off)}}(a)=\mathbb{E}_{R(a)\sim P^{\text{(off)}}_{a}}[R(a)]$. We do not have any boundedness assumption on $\mu^{\text{(on)}}(a),\mu^{\text{(off)}}(a)$.
@@
$\text{Reg}_{T}(\pi,P)=T\mu_{*}^{\text{(on)}}-\sum_{t=1}^{T}\mu^{\text{(on)}}(A_{t})=\sum^{T}_{t=1}\Delta(A_{t}),$ (3)
@@ 2.1  Auxiliary Input: Valid Bias Bounds $V$
We say that $V=\{V(a)\}_{a\in{K}}\in(\mathbb{R}_{\geq 0}\cup\{\infty\})^{|\mathcal{A}|}$ is a valid bias bound on an instance $I$ if
$V(a)\geq|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|\qquad\text{for each $a\in{\cal A}$.}$ (4)
@@
In our forthcoming policy MIN-UCB, the DM is endowed the auxiliary input $V$ in addition to the offline dataset $S$ before the online phase begins. The quantity $V(a)$ serves as an upper bound on the amount of distributional shift from $P^{\text{(off)}}_{a}$ to $P^{\text{(on)}}_{a}$. The condition (4) always holds in the trivial case of $V(a)=\infty$ for all $a\in{\cal A}$, which is when the DM has no knowledge on the relationship between $P^{\text{on}},P^{\text{off}}$. By contrast, when $V(a)\neq\infty$, the DM has non-trivial knowledge on the difference $|\mu^{\text{(off)}}(a)-\mu^{\text{(on)}}(a)|$.
@@
The knowledge of an upper bound such as $V$ is in line with the model assumptions in studies on learning under distributional shift. Similar upper bounds are assumed in the contexts of supervised learning [CrammerKW08], stochastic optimization [besbesMM22], offline policy learning for contextual bandits [SiNZB23], and multi-task bandit learning [WangZSRC21]. Methodologies for constructing such bounds are also discussed in the literature. For instance, blanchet2019robust designs an approach based on several machine learning estimators, such as LASSO. SiNZB23 provides managerial insight on how to estimate such bounds empirically. chen2022data constructs them through cross-validation and implements it in a healthcare setting.
2.2  Application: Dynamic Pricing
Now we discuss the application of dynamic pricing (keskin2014dynamic), which is naturally captured by our model and helps explain the rationale behind introducing offline data.
In the single-product dynamic pricing setup, a retailer dynamically adjusts the posted price of a product for sequentially arriving customers, aiming to maximize total revenue. Specifically, the arm set $\mathcal{A}$ consists of $K$ possible prices. Each customer can either purchase one unit of the product or do nothing. The customer utility values for the product are i.i.d. according to an unknown latent probability distribution $\Gamma$. At each time $t$, one customer arrives, and the DM posts a price $A_{t}\in\mathcal{A}$. The customer makes a purchase if and only if her/his random utility $U$ is larger than or equal to $A_{t}$, i.e., $U\geq A_{t}$. In which case the customer purchases, she/he will pay $A_{t}$ to the DM. Thus, $R_{t}=\boldsymbol{1}\{U\geq A_{t}\}\cdot A_{t}$ and $\boldsymbol{\mu}^{\text{(on)}}(a)=a\cdot\Pr(U\geq a)$. The DM, however, lacks prior information about the latent utility distribution $\Gamma$ and only observes binary purchase feedback at the posted price.
In many practical scenarios, the DM has access to offline data, past sales observations from related products or different markets. Such data can provide valuable information for learning the demand function and guiding future pricing, even if not perfectly aligned with the current process. For example, when a retailer plans to sell a product to a new customer group that overlaps substantially (e.g. 90%) with a previously served group, historical pricing data naturally inform new pricing strategies. Since only a small fraction of customers differ, the aggregrated price sensitivity of the new group is expected to be similar to that of the previous one, i.e. $P^{\text{(off)}}$ and $P^{\text{(on)}}$ are close in our model. It is therefore reasonable to assume that a valid bias bound $V$ with small $V_{\max}$ exists and can be estimated via machine learning or statistic tools, as discussed previously.
The relevance of such offline data is further reflected in many real-world scenarios. For example, when a company launches a product in Canada after its release in the United States, the two markets typically share substantial similarities, such as language, culture. Although there exist certain differences, such as variations in income levels and regional regulations, these differences are relatively minor, and the U.S. sales data can still provide valuable insights for pricing decisions in the Canadian market. Likewise, when Apple releases a new-generation iPhone, like iPhone 17, the target customers largely overlap with those of the previous model, such as loyal users in Apple ecosystem. While the new model may differ in some features or technology, the customer preferences and price sensitivity remain broadly consistent. In both examples, offline data not only exist but also have tangible modeling and algorithmic value, that they can meaningfully accelerate learning and enhance pricing performance in practice.
@@ 3  An Impossibility Result
We illustrate our impossibility result with two instances $I_{P},I_{Q}$. The instances $I_{P},I_{Q}$ share the same arm set ${\cal A}=\{1,2\},\{T_{\text{S}}(a)\}_{a\in{\cal A}}$ and horizon length $T$. However, $I_{P},I_{Q}$ differ in their respective reward distributions $P$, $Q$, and $I_{P},I_{Q}$ have different optimal arms. Consider a fixed but arbitrary $\beta\in(0,1/2)$. Instance $I_{P}$ has well-aligned reward distributions $P^{\text{(off)}}=P^{\text{(on)}}$, where
$P^{\text{(on)}}_{1}={\cal N}(0,1),\qquad P^{\text{(on)}}_{2}={\cal N}(-T^{-\beta},1).$
In instance $I_{P}$, we have $\Delta(2)=T^{-\beta}$, and a non-empty offline dataset provides useful hints on arm 1 being optimal. For example, given $T_{\text{S}}(1)=T_{\text{S}}(2)\geq 128(T^{2\beta}-T^{2\beta-\epsilon})\log T$ where $\epsilon\in(0,\beta)$, the existing policies H-UCB and MonUCB achieve an expected regret (see Equation (2)) at most $O\left(\frac{\log T}{\Delta(2)}-T_{\text{S}}(2)\Delta(2)\right)=O(\log T(T^{\beta}-T^{\beta}+T^{\beta-\epsilon}))=O(T^{\beta-\epsilon}\log T)$, which strictly outperforms the vanilla UCB policy that incurs expected regret $O(\log T/\Delta(2))=O(T^{\beta}\log T)$. Despite the apparent improvement, the following claim shows that any non-anticipatory policy (not just H-UCB or MonUCB) that outperforms the vanilla UCB on $I_{P}$ would incur a larger regret than the vanilla UCB on a suitably chosen $I_{Q}$.
Theorem 1.
Let $T_{\text{S}}(1),T_{\text{S}}(2)$ be arbitrary. Consider an arbitrary non-anticipatory policy $\pi$ (which only possesses the offline dataset $S$ but not the auxiliary input $V$) satisfies $\mathbb{E}[\text{Reg}_{T}(\pi,P)]\leq CT^{\beta-\epsilon}\log T$ on instance $I_{P}$, where $\epsilon\in(0,\beta)$, $C>0$ is an absolute constant, and the horizon $T$ is so large that $C<T^{\epsilon}/(4\log T)$. Set $Q=(Q^{\text{(off)}},Q^{\text{(on)}})$ in the instance $I_{Q}$ as $Q^{\text{(off)}}=P^{\text{(off)}}$, $Q^{\text{(on)}}_{1}=P^{\text{(on)}}_{1}$,
@@
The following inequality holds: $\mathbb{E}[\text{Reg}(\pi,Q)]\geq$
$\frac{T^{1-\beta}}{4}\cdot e^{-2}-CT^{\beta-\epsilon}\log T=\Omega(T^{1-\beta})>\Omega(\sqrt{T}).$ (5)
@@
We first remark on $I_{Q}$. Different from $I_{P}$, arm 2 is the optimal arm in $I_{Q}$. Thus, the offline data from $Q^{\text{(off)}}$ provides the misleading information that arm 1 has a higher expected reward. In addition, note that $Q^{\text{(off)}}=P^{\text{(off)}}$, so the offline datasets in instances $I_{P},I_{Q}$ are identically distributed. The theorem suggests that no non-anticipatory policy can simultaneously (i) establish that $S$ is useful for online learning in $I_{P}$ to improve upon the UCB policy, (ii) establish that $S$ is to be disregarded for online learning in $I_{Q}$ to match the UCB policy, negating the short-lived conjecture in (Q).
@@
Lastly, we allow $T_{\text{S}}(1),T_{\text{S}}(2)$ to be arbitrary. In the most informative case of $T_{\text{S}}(1)=T_{\text{S}}(2)=\infty$, meaning that the DM knows $P^{\text{(off)}},Q^{\text{(off)}}$ in $I_{P},I_{Q}$ respectively, the DM still cannot achieve (i, ii) simultaneously. Indeed, the assumed improvement $\mathbb{E}[\text{Reg}_{T}(\pi,P)]\leq CT^{\beta-\epsilon}\log T$ limits the number of pulls on arm 2 in $I_{P}$ during the online phase. We harness this property to construct an upper bound on the KL divergence between on online dynamics on $I_{P}$, $I_{Q}$, which in turns implies the DM cannot differentiate between $I_{P},I_{Q}$ under policy $\pi$. The KL divergence argument utilizes a chain rule (see Appendix D.1) on KL divergence adapted to our setting.
3.1  Proof for Theorem 1
We first set up some notation. Consider a non-anticipatory policy $\pi$ and an instance $I$ with reward distribution $P$. We denote $\rho_{P,\pi}$ as the joint probability distribution function on $S,A_{1},R_{1},\ldots,A_{T},R_{T}$, the concatenation of the offline dataset $S$ and the online trajectory under policy $\pi$ on instance $I$. For a $\sigma(S,A_{1},R_{1},\ldots,A_{T},R_{T})$-measurable event $E$, we denote $\Pr_{P,\pi}(E)$ as the probability of $E$ holds under $\rho_{P,\pi}$. For a $\sigma(S,A_{1},R_{1},\ldots,A_{T},R_{T})$-measurable random variable $Y$, we denote $\mathbb{E}_{P,\pi}[Y]$ as the expectation of $Y$ under the joint probability distribution function $\rho_{P,\pi}$. In our analysis, we make use of $Y$ being $\text{Reg}_{T}(\pi,P)$ or $N_{T}(a)$ for an arm $a$. To lighten our notational burden, we abbreviate $\mathbb{E}_{P,\pi}[\text{Reg}_{T}(\pi,P)]$ as $\mathbb{E}[\text{Reg}_{T}(\pi,P)]$. By the condition $C<T^{\epsilon}/(4\log T)$, we know that
$\frac{1}{\sqrt{C\log T}T^{\beta-(\epsilon/2)}}-\frac{1}{T^{\beta}}>\frac{1}{T^{\beta}}>0,$
so in the instance with $Q$, arm 2 is the unqiue optimal arm. Consider the event $E=\{N_{T}(2)>T/2\}$. Then $\mathbb{E}[\text{Reg}(\pi,P)]+\mathbb{E}[\text{Reg}(\pi,Q)]$
$\geq$ $\frac{1}{T^{\beta}}\cdot\frac{T}{2}\cdot\Pr_{P,\pi}(E)$
$+\left[\frac{1}{\sqrt{C\log T}T^{\beta-(\epsilon/2)}}-\frac{1}{T^{\beta}}\right]\cdot\frac{T}{2}\Pr_{Q,\pi}(E^{C})$
$\geq$ $\frac{T^{1-\beta}}{4}\cdot\exp\left[-\mathbb{E}_{\pi,P}[N_{T}(2)]\cdot\text{KL}(P^{(\text{on})}_{2},Q^{(\text{on})}_{2})\right$.
$\left.-T_{S}(2)\cdot\text{KL}(P^{(\text{off})}_{2},Q^{(\text{off})}_{2})\right$] (6a)
$=$ $\frac{T^{1-\beta}}{4}\cdot\exp\left[-\mathbb{E}_{\pi,P}[N_{T}(2)]\cdot\text{KL}(P^{(\text{on})}_{2},Q^{(\text{on})}_{2})\right$]. (6b)
(6a) is again by the chain rule in Theorem 11 in the Appendix. (6b) is because $P^{\text{(off)}}=Q^{\text{(off)}}$. It is evident that $\text{KL}(P^{(\text{on})}_{2},Q^{(\text{on})}_{2})=1/(2CT^{2\beta-\epsilon}\log T)$, and by the Claim assumption that $\text{Reg}(\pi,P)=(1/T^{\beta})\mathbb{E}_{\pi,P}[N_{T}(2)]\leq CT^{\beta-\epsilon}\log T$, we have $\mathbb{E}_{\pi,P}[N_{T}(2)]\leq CT^{2\beta-\epsilon}\log T$, which leads to
Again by the Theorem assumption $\mathbb{E}[\text{Reg}(\pi,P)]\leq CT^{\beta-\epsilon}\log T$, we arrive at the claimed inequality, and the Theorem is proved.
4  MIN-UCB: Leverage Offline Data with Auxiliary Information